### v1

# SMC-X: A Distributed Scalable Monte Carlo Simulation Method for Chemically Complex Alloys

Abstract
To predict the complex chemical evolution in multicomponent alloys, it is highly desirable to have accurate atomistic simulation methods capable of reaching sufficiently large spatial and temporal scales. In this work, we advance the recently proposed SMC-X method through distributed computation on either GPUs or CPUs, pushing both spatial and temporal scales of atomistic simulation of chemically complex alloys to previously inaccessible scales. This includes a 16-billion-atom HEA system extending to the micrometer regime in space, and a 1-billion-atom HEA evolved over more than three million Monte Carlo swap steps, approaching the minute regime in time. We show that such large-scale simulations are essential for bridging the gap between experimental observations and theoretical predictions of the nanoprecipitate sizes in HEAs, based on analysis using the Lifshitz–Slyozov–Wagner (LSW) theory for diffusion-controlled coarsening. This work demonstrates the great potential of SMC-X for simulation-driven exploration of the chemical complexity in high-entropy materials at large spatial and temporal scales.
keywords:
American Chemical Society,
[…]
{tocentry}
1 Introduction
Materials are composed of atoms, and the types of atoms together with their spatial arrangements govern a wide spectrum of physical properties, including mechanical, electrical, magnetic, and thermal behavior. However, the vast chemical complexity present in many technologically important materials poses formidable challenges for existing computational approaches. This challenge is particularly pronounced in high-entropy alloys (HEAs) 1, which contain multiple principal elements and therefore exhibit substantially increased chemical complexity. Moreover, this complexity is further compounded by the interplay and competition between order and disorder, as well as enthalpic and entropic effects, across diverse spatial, temporal, and temperature scales, making predictive modeling of such systems exceptionally demanding. Capturing these complexities, including chemical composition 2, disorder 3, chemical short-range order (SRO) 4, 5, chemical medium-range order (MRO) 6, multi-component intermetallic nanoparticles (MCINP) 7, and nano lamellae8, 9, requires modeling and simulating atomic systems over large spatial and temporal scales, with accuracy and generalizability of first principles method 10, 11, 12.
In recent years, machine learning potentials (MLPs) 13, 14, 15 trained on DFT datasets 16, 17 have emerged as a highly promising solution to the aforementioned bottlenecks 18, 19. By fitting interatomic potential energy surfaces with machine learning models, MLPs can retain near-DFT accuracy 20 while dramatically increasing simulation efficiency and extending accessible system sizes 21. With an appropriate interaction cut-off radius, many MLPs can reduce the computational complexity of traditional DFT calculations from cubic scaling, $O(N^{3})$, to linear scaling, $O(N)$ 15, with respect to the number of atoms $N$. Moreover, by bypassing the explicit solution of the electronic structure, the computational prefactor can be reduced by orders of magnitude 22, enabling simulations at spatial scales previously deemed intractable 23. The integration of MLPs with molecular dynamics (MD), a foundational atomistic simulation technique renowned for its excellent scaling, has revolutionized computational materials science 24, chemistry 25, and biology 26 by enabling accurate, large-scale simulations at significantly reduced cost.
Although the MLP + MD approach successfully scales simulations to unprecedented sizes 27, 28, 29, its intrinsic limitations restrict accessible time scales to the nanosecond regime, which typically requires millions of sequential MD steps. Consequently, it is extremely challenging, if not impossible, for MD to capture many important slow-evolving phenomena, such as precipitate formation in alloys 7. For instance, a single swap step between two neighboring lattice sites may take millions or even billions of MD steps to simulate. For such a purpose, Monte Carlo (MC) simulations offer unique advantages due to their characteristic of directly sampling configurational space, rather than tracking real-time trajectories as in MD 30, 31. However, compared to MD, the integration of conventional MC with MLP faces difficulties when scaled up to large systems 22. This challenge stems from the intrinsically sequential update nature in mainstream Monte Carlo algorithms, such as the Metropolis method, where trial moves are proposed site by site, inherently limiting parallelization 32. This bottleneck is in stark contrast with MD, where all atoms can be updated simultaneously at each step.
Previous efforts to scale Monte Carlo (MC) simulations have introduced parallel trial moves using generalized checkerboard algorithms, with most implementations customized for specific interaction types or hardwares 33, 34, 35, 36, 37, 32, and only a few designed for general interactions 38, 22. Among all these implementations, the recently proposed scalable Monte Carlo at eXtreme (SMC-X) 22 method stands out due to its superb efficiency. Such a high efficiency is achieved via exposing the previously hidden degree of parallelism by dynamical link-cell (LC) decomposition and local interaction zone (LIZ), and by carefully mapping these degrees of parallelisms onto modern GPU architectures to remove unnecessary data movement. These innovations allow SMC-X to simulate systems as large as one billion atoms using a single GPU 22. This performance not only substantially exceeds all prior MC + MLP methods, but also exhibits advantages against state-of-the-art (SOTA) MD + MLP implementations in both the number of atoms and the atomistic simulation throughput, measured by $N_{\rm atom}\times N_{\rm steps}$ per second 22. A comparison of various computational methods for chemically complex alloys is shown in Tab. 1. It should be noted that while mesoscale methods such as the Potts model, kinetic Monte Carlo, and phase-fild methods can simulate very large spatial and temporal scales, they introduce coarse-grained approximations that rely on empirical parameters, making them less rigorous at the atomistic level, therefore are to compared in Tab. 1.
[…]
Method Length Time Primary Strength/ Limit Typical Applications
DFT Å–nm ps High accuracy and generalizability, but computationally expensive Phonon dispersion, stacking fault energies
[…]
MLP + Hybrid MC/MD nm ms Can handle both structural relaxation and chemical complexity, but lost true kinetics. chemical SRO, GB segregation
[…]
In this work, we advance the SMC-X method to even larger spatial and temporal scales through distributed computing. We demonstrate that, by carefully mapping the algorithmic parallelism in SMC-X into the hardware parallelism in modern computational systems, we can achieve atom-by-atom simulation of chemically complex alloys at unprecedented spatial and temporal scales. We further demonstrate the versatility of SMC-X across different computation platforms with the efficient implementations on both the CPUs and GPUs. Collectively, these advances establish SMC-X as a framework that pushes the spatial and temporal scales of atomistic simulation to unparalleled levels, providing a crucial bridge between experiment and theory, as exemplified by the prediction of nanoprecipitate size in high-entropy alloy. An illustration of the method in this work is shown in Fig. 1.
Figure 1: A schematic of the workflow to illustrate the main contributions. MLP: machine learning potential; LC: link-cell; LIZ: local interaction zone; T: temperature; SIMD: single-instruction multiple-data; $S$: solid solution; SRO: short-range order; LRO: long-range order; MCINP: multicomponent intermetallic nanoparticle.
2 Theory and Methods
In this section, we present an introduction of the theoretical background of distributed SMC-X. Note that the basic concept of the SMC-X algorithm has been introduced in Ref. 22, but the focus was on its application in materials science. Here we present a more detailed discussion and introduce the distributed implementation on both CPUs and GPUs.
2.1 Monte Carlo simulation
In MC, physical observables are computed as ensemble averages over sampled configurations, providing direct access to thermodynamic properties. For instance, in the canonical ensemble (NVT), the probability of observing a configuration $\bm{\sigma}_{k}$ is given by:
$P(\bm{\sigma}_{k})=\frac{e^{-\beta H(\bm{\sigma}_{k})}}{\sum_{\bm{\sigma}_{k}}e^{-\beta H(\bm{\sigma}_{k})}},$ (1)
where $\beta=1/(k_{B}T)$, $k_{B}$ is the Boltzmann constant, and $T$ is the temperature. Physical observables $\langle\hat{O}\rangle$ such as the expected energy, magnetization, specific heat, and magnetic susceptibility can then be calculated accordingly from
$\langle\hat{O}\rangle=\sum_{k}\frac{O(\sigma_{k})e^{-\beta H(\bm{\sigma}_{k})}}{Z}.$ (2)
A sample following the probability distribution of Eq. 1 can be generated efficiently via Markov chain Monte Carlo (MCMC) simulation. For such a purpose, the Metropolis-Hastings algorithms is the most commonly used method. The Metropolis-Hasting algorithm starts with proposing a move (trial) at each lattice site, with the acceptance probability of this trial given by:
$P=\begin{cases}1,&\Delta E\leq 0\\ \exp(-\beta{\Delta E}),&\Delta E>0.\end{cases}$ (3)
A Monte Carlo step, or sweep, is defined as attempting exactly one trial move per lattice site. An MC simulation typically begins with a series of warm‑up steps to ensure thermal equilibrium. After equilibration, the configurations are measured in subsequent steps, allowing the recorded samples to be used to estimate physical observables via statistics.
The trial moves within the Metropolis-Hasting algorithm are inherently sequential, which poses a significant challenge to parallelization. Although advanced techniques such as cluster MC 39 or the Wang–Landau algorithms 40, 41, 42 can alleviate this limitation by reducing or avoiding sequential site-by-site updating, they are often restricted in scope and lack general applicability. A more general approach leverages the assumption of short-range interactions: If interactions are sufficiently localized, then spatially distant regions can be updated independently and in parallel. In the simple Ising model, this principle underlies the well-known checkerboard algorithm. Extensions of this approach to three-dimensional lattices and beyond nearest-neighbor interactions are conceptually straightforward for specific models 33. However, generalization to arbitrary short-range interactions is more limited. One notable example is the Scalable Parallel Monte Carlo (SPMC) algorithm 38, as implemented in the LAMMPS package 43. Despite its utility, SPMC has disadvantage for further scaling up: to minimize the weak spatial correlations in SPMC algorithm, the total volume of the domains need to be maximized 38, which underutilizes the available parallelism, especially on architectures with a high degree of concurrency. As a result, current application of SPMC are limited systems of less than a million atoms 29, 44.
Figure 2: A schematic to illustrate the SMC-X algorithm, including link-cell (LC), local interaction zone (LIZ), and domain decomposition scheme, using the 2D square lattice as an example. LIZ+ represents sites needed for calculating the energy change due to possible MC trial. Adapted from Ref.  22
2.2 Distributed SMC-X
In the SMC-X framework 22, the challenge of site-by-site sequential updating is addressed through two key innovations: dynamical link-cell (LC) and local interaction zone (LIZ), as illustrated in Fig. 2. Compared to SPMC 38, the LCs are dynamically constructed to ensure that every site is visited exactly once per MC step, and that the detailed balance condition is strictly satisfied. The resulting degree of parallelism is given by:
$P_{LC}=N_{atom}/N_{LC},$ (4)
where $N_{atom}$ is the total number of atoms in the supercell, and $N_{LC}$ is the number of atoms within LC. By choosing the smallest LC size that fully separates short-range interactions, the parallelism $P_{\mathrm{LC}}$ can be maximized. The LIZ is introduced to harness the short-range interaction to reduce the computational complexity of a MC trial move from $O(N)$ to $O(1)$. Furthermore, by computing the atomic features within each LIZ on the fly and tightly couple it with the MC simulation, the memory usage in SMC-X can be drastically reduced, and near-core caches can be effectively utilized to significantly enhance the computational throughput.
In addition to LC and LIZ, there are additional parallel opportunities in SMC-X, as illustrated in the algorithm part in Fig. 1. Among them, the lattice and the temperature parallelism can be implemented as MPI ranks for distributed computation over multiple chips. In lattice parallelism, the complete lattice is partitioned across multiple GPUs, with a ghost layer used to exchange boundary information between neighboring sublattices, as illustrated in Fig. 2. Under this scheme, if the number of atoms per GPU remains fixed, the total system size then increases proportionally to the number of GPUs, demonstrating weak scaling. Conversely, when the same system is distributed over more GPUs, the time required for a single MC step decreases, enabling more steps to be simulated within the same wall-clock time, an indication of strong scaling. Therefore, the strong and weak scaling behaviors in lattice parallelism represent the potential to push the atomistic simulation to larger spatial and temporal scales, respectively, by simply adding more computational resources.
The temperature parallelism can be discussed in two different situations. First, note that in the canonical ensemble, temperature is fixed, and simulations at different temperatures are essentially independent. Therefore, the MC simulation is embarrassingly parallel when properties at different temperatures are needed, such as to search for an order-disorder transition temperature. On the other hand, it is well known that the convergence speed of conventional MC can be slow, particularly near second-order transition, and the replica-exchange (RE, also known as parallel tempering) is a widely used technique to address this problem. In RE, multiple replicas of the same system are simulated concurrently at different temperatures. Configurations between replicas are periodically exchanged according to a Metropolis criterion that preserves detailed balance. The transition probability for two configurations with energies $E_{i}$ and $E_{j}$ is given by:
$P_{\text{swap}}=\rm{min}(1,\exp(\Delta)),$ (5)
where
$\Delta=\frac{1}{k_{B}}\left(\frac{1}{T_{i}}-\frac{1}{T_{j}}\right)(E_{i}-E_{j}),$ (6)
$T_{i}$ and $T_{j}$ are the temperatures for configuration $i$ and $j$. By exchanging configurations between neighboring temperatures, the replica-exchange algorithm mitigates the risk of the system becoming trapped in local energy minima. This is especially critical near second-order phase transition points 39, 40, where the divergence of the correlation length makes it challenging for simulations at a fixed temperature to escape metastable states. For large systems, an additional complication needs to be considered: the energy fluctuation decreases as the system size increases. Consequently, the energy distributions of adjacent replicas may overlap insufficiently, leading to low exchange rates. To ensure reasonable exchange rate, the temperature difference $\Delta T$ need to be adjusted in accordance with the system size $N_{\text{atom}}$ as:
$\Delta T\propto 1/\sqrt{N_{\text{atom}}},$ (7)
For large atomic systems, the required temperature spacing $\Delta T$ becomes very small (e.g., 0.1 K for 1 billion atoms), which significantly increases the number of replicas and thus the overall computational cost. Limited by computing resources, we do not attempt a full RE simulation in this work. Note that the $\rm{Fe_{29}Co_{29}Ni_{28}Al_{7}Ti_{7}}$ HEA studied in this work does not show obvious order-disorder transition in its $C_{v}$ curve, as illustrated in Ref. 22.
2.3 SMC-GPU and SMC-CPU
As shown in Fig. 1, the SMC-X framework contains a hierarchy of algorithmic parallelisms. To efficiently utilize these opportunities, the various algorithmic parallelism need to be carefully mapped into the hardware parallelism in modern high-performance computers, in which the main computing capability typically coming from CPUs or GPUs. Considering the differences in CPU and GPU, such as hardware architecture and programming model, we designed different variants of SMC-X: SMC-GPU and SMC-CPU, as illustrated in Fig. 3.
On CPUs, fine-grained tasks such as ML model evaluations and local energy updates are accelerated by vector and matrix units, while medium-grained tasks such as link-cell and lattice decomposition are distributed over CPU cores and intranode processors using SIMD, OpenMP, and MPI programming. Coarse-grained tasks, such as temperature parallelism via replica exchange, are mapped across computing nodes. On GPUs, an analogous hierarchy is employed: CUDA and tensor cores handle fine-grained ML operations, streaming processors execute local interaction and link-cell computations, and intranode GPUs coordinate lattice decomposition. Finally, replica exchange across temperatures is distributed over GPU computing nodes via CUDA/MPI. This hierarchical mapping strategy ensures that algorithmic parallelism is effectively matched to the granularity of available hardware parallelism, thereby maximizing scalability and efficiency of SMC-X simulations.
Figure 3: Illustration of the SMC-GPU and SMC-CPU algorithms. In SMC-GPU, the LIZ and LC degree of parallelism are allocated to the cuda.blocks and cuda.threads. In SMC-CPU, part of the LC degree of parallelism (x and y directions) are parallelized via MPI processes, while the other LC degree of parallelism, as well as the LIZ ones, are parallelized via OpenMP threads.
3 Results and discussion
In this section, we present the simulation results for the $\rm{Fe_{29}Co_{29}Ni_{28}Al_{7}Ti_{7}}$ HEA. Two models, EPI and qSRO, are used. The EPI model31, 45 contains only pairwise interactions, which is computationally highly efficient since there is no need to explicitly evaluate the local energy changes within the LIZ. On the other hand, the qSRO model 22 represents a general nonlinear short-range model, where the local energies of all lattice sites within LIZ are calculated to determine the energy change due to a MC move. For SMC-GPU, the computational platform is a 2-node GPU cluster. Each computational node contains 8 80-GB NVIDIA H800 GPUs interconnected via 400 GB/s NVLink, and the 2 nodes are connected with a 200 GB/s InfiniBand network. For SMC-CPU, the platform is a node with 2 Kungpeng ARM CPUs. Each CPU contains multiple NUMA domains, with dozens of physical cores in each domain.
The DFT dataset is calculated with the LSMS method 46. LSMS is an all-electron electronic structure calculation method in which the computational cost scales linearly with respect to the number of atoms. For FeCoNiAlTi, we use a 100 atom supercell with a lattice constant of 6.72 Bohr (0.356 nm). We employ a spin-polarized scheme to account for the magnetic interactions in the system. The angular momentum cutoff $l_{max}$ for the electron wavefunctions is chosen as 3, and the LIZ is chosen as 86 atoms. We used PBE as the exchange-correlational functional. To enhance the representativeness of the DFT data, the total dataset is made up of 120 randomly generated configurations and 100 configurations from Monte Carlo simulation, as proposed in 31. Therefore, the total number of configurations is 220, with each comprising 100 atoms.
3.1 Performance
Figure 4: Comparison of different ML accelerated atomistic simulation methods in terms of atom per chip and throughput per chip.
Table 2: Comparison of the performance of atomistic simulations achieving DFT-level accuracy (10 meV/atom). Throughput (TP) is quantified as $N_{\text{atom}}\times N_{\text{step}}/T$, where $T$ is the simulation time in seconds, with $\text{atom}\cdot\text{step}/\text{s}$ as the unit. DP stands for DeepPotential.
Simu. Potential Sys. Chips Machine Atoms TP TP/Chip
LAMMPS 27 Allegro HIV 5,120 Perlmutter 100 M $1.27\times 10^{10}$ $2.47\times 10^{6}$}
LAMMPS 47 DP H2O 90,000 Sunway 29 B $1.14\times 10^{10}$} $1.26\times 10^{5}$
LAMMPS 48 TensorMD W 32,000 Sunway 51.8 B $\mathbf{{1.30\times 10^{10}}}$ $4.06\times 10^{5}$
GPUMD 49 UNEP-v1 Cu 8 A100 1 node 100 M $1.50\times 10^{8}$} $1.86\times 10^{7}$
GPUMD 50 NEP89 HEA 1 H800 1 node 15 M $7.0\times 10^{6}$} $7.0\times 10^{6}$
SMC-X EPI HEA 16 H800 2-node 16 B $\mathbf{5.63\times 10^{10}}$ $\mathbf{3.52\times 10^{9}}$
SMC-X qSRO HEA 16 H800 2-node 16 B $5.79\times 10^{8}$ $\mathbf{3.62\times 10^{7}}$}
SMC-X qSRO HEA 2 CPU 1 node 16 B $4.44\times 10^{7}$ $2.22\times 10^{7}$}}
The performance is measured with three metrics: number of atoms, which measure the spatial scale; simulation throughput, which measures time to solution; and scaling efficiency, which measures the scale-up potential with increasing computational resources. As shown in Fig. 4, the EPI model exhibits exceptional computational efficiency on a single GPU, achieving a throughput of $3.5\times 10^{9}$ $\text{atom}\cdot\text{step}/\text{s}$—two orders of magnitude faster than the qSRO_GPU result of $3.7\times 10^{7}$ $\text{atom}\cdot\text{step}/\text{s}$. In comparison, the single-NUMA performance of qSRO_CPU reaches $3.5\times 10^{6}$ $\text{atom}\cdot\text{step}/\text{s}$, or $2.5\times 10^{7}$ $\text{atom}\cdot\text{step}/\text{s}$ per CPU. The throughput of SMC-X not only far exceeds that of traditional Monte Carlo methods but also outperforms state-of-the-art molecular dynamics implementations, as shown in Fig. 4 and summarized in Tab. 2. The advantage is even more pronounced in terms of the number of atoms supported on a single chip: the SMC-GPU implementation accommodates up to 1 billion atoms on a single GPU, whereas the best reported MD simulation to date reaches only 15 million atoms per GPU, as shown in the row of NEP89 in Tab. 2. As mentioned, these performance gains stem from key implementation optimizations, including on-the-fly computation and tight coupling between the atomistic simulation and machine learning models, which collectively minimize data movement and maximize computational efficiency.
Figure 5: The strong and weak scaling of SMC-GPU and SMC-CPU with EPI and qSRO model. CCO represents computation-communication-overlap.
The strong and weak scaling efficiencies of various methods are shown in Fig. 5. We also implemented a computing-communication overlap for EPI_GPU, as efforts to enhance the scaling efficiency for communication heavy cases. The results of all the above cases are shown in Fig. 5. First, we find that the qSRO model consistently demonstrates superior scaling efficiency due to its higher computation-to-communication ratio. For example, the strong scaling efficiency of qSRO_GPU reaches 0.934 on 16 GPUs, compared to only 0.339 for EPI_GPU. Second, the strong scaling of EPI_GPU can be substantially improved from 0.339 to 0.679 on 16 GPUs by applying computation–communication overlap (CCO). Notably, on 8 GPUs, CCO enhances efficiency from 0.713 to 0.897. These results highlight that CCO is highly effective in improving scaling behavior, particularly when combined with high-speed interconnects (e.g., Nvidia NVLINK within 8 GPUs) and large system sizes. Third, all methods exhibit excellent weak scaling, with 16-GPU efficiencies of 0.915, 1.007, 0.985, and 0.791 for EPI_GPU, EPI_GPU_CCO, qSRO_GPU, and qSRO_CPU, respectively. This demonstrates that SMC-X can be readily extended to larger spatial scales. Finally, although both strong and weak scaling efficiencies for qSRO_CPU are generally good, they are consistently lower than their GPU counterparts, indicating that cross-CPU NUMA communication is more costly than GPU interconnects such as NVLINK.
3.2 Evolution of chemical defects
The Warren–Cowley short-range order (SRO) parameter 51, 31, denoted as $\alpha_{m}^{pp^{\prime}}$, is widely used to quantify local chemical ordering. It is defined as
$\alpha_{m}^{pp^{\prime}}=1-\frac{P^{p|p^{\prime}}_{m}}{c_{p}},$ (8)
where $c_{p}$ is the concentration of species $p$, and $P^{p|p^{\prime}}_{m}$ is the conditional probability of finding an atom of type $p$ in the $m$-th neighbor shell around a central atom of type $p^{\prime}$. The SRO parameters are obtained by averaging over the Monte Carlo configurations. By definition, $\alpha=0$ corresponds to perfect randomness, $\alpha<0$ indicates a tendency toward ordering between $p$ and $p^{\prime}$, and $\alpha>0$ reflects a tendency toward segregation. Figure 6 shows the nearest-neighbor (NN) SRO parameters obtained from Monte Carlo simulations of a one-million-atom FeCoNiAlTi system. At high temperature, the alloy remains nearly random, whereas at lower temperatures distinct SRO develops. In particular, Ti and Ni atoms display a strong tendency to segregate from Fe and Co, with the Ti–Fe pair showing the strongest separation. This behavior is consistent with previous experimental 7 and theoretical 22 studies, which reported the formation of $L1_{2}$-type nanoprecipitates enriched in Ni and Ti within a Fe- and Co-rich matrix.
Figure 6: The short range order parameters calculated with the 1 million atom supercell. (a) The SROs for different elements pairs; (b) The SROs for the same element pairs.
To investigate microstructure evolution, we performed large-scale Monte Carlo simulations using the EPI model, with a supercell comprising 1,066,867,200 atoms (dimensions: $672\times 630\times 630\times 4$). The computing platform is the two node, 16 H800 GPU cluster, and the ML model is the EPI model to maximize the simulation speed. We decrease the temperature from 2000 K to 1000 K, with tens of thousands of MC steps at each temperature, except for a temperature we refer to as the “aging” temperature $T_{aging}$, at which the simulation run for 3 million MC steps . The X-Y face of the system for $T_{aging}=1000$ K experiment is shown in Fig. 7 (a, b), where we used 8 GPUs for lattice parallelism and 2 temperature degree of parallelism, with $\Delta T=0.06$ K. As mentioned in the method section, such a small $\Delta T$ is for reasonable acceptance ratio of RE in large supercells. The results for $T_{aging}=1500$ K is shown in Fig. 7 (c, d), where no RE is used, and the lattice is parallelized over the total 16 GPUs. From these figures, it can be seen that the NPs rich in Ni and Ti already forms at 1500 K. Upon further cooling, nanoprecipitates gradually grow in size. Moreover, the comparison between Fig. 7 (a) and (c) suggests that the two temperature replicas demonstrate very little effect, which is consistent with our expectation, due to the tiny value of $\Delta T$. Finally, it can be seen that the size of NPs increases as $T_{aging}$ switched from 1000 K to 1500 K, indicating the high rate of NP formation at elevated temperature.
Figure 7: Simulation results at different temperature schedule for 1-billion-atom $\rm{Fe_{29}Co_{29}Ni_{28}Al_{7}Ti_{7}}$. Note that two FCC layers are shown in a checkerboard grid for clarity of the image. (a-d) The X-Y faces at different stages of the MC simulation, as denoted in the subtitles. Atoms are colored as Fe: red; Co: purple; Ni: green; Al: grey; Ti: blue.
3.3 APT sample
Figure 8: Comparison of APT sample from the simulation with previous experimental and theoretic studies. (a) The APT sample extracted from SMC-X simulation in Ref. 22. (b) The APT reconstructed sample image of the same materials from experiment 7, where MCM represents multicomponent matrix, and MCINP represents multicomponent intermetallic nanoparticle. (c) The APT sample from this work, using 3 million MC steps at 1000 K. (d) The APT sample extracted from Fig. 7 d, using 3 million MC steps at 1500 K. (a-c) blue represents MCINP, and red represents MCM.
To reveal further details of the NPs, we extract an atom-probe-tomography (APT) sample from the simulation results and compared them with previous experimental and theoretical studies, as shown in Fig. 8. First, we see that the previous study shown in Fig. 8(a) clearly underestimates the sizes of the NPs, which is about 7 nm in diameter, as compared to the experimental result in Fig. 8 (b), which is about 30 nm. Second, we see that this gap between experiment and theory can be reduced by increasing the MC steps from 0.1 million to 3 million, which enhances the size of the NP to about 14 nm. Finally, note that the experimental sample is synthesized by homogenizing at 1423 K for 2 h and aged at 1053 K for 4 h. The 3 million steps at 1500 K in Fig. 8(d) can be seen as an effort to mimic the homogenization effect. It can be seen that the size of the NP can be further increased to about 20 nm. These observations reveal that larger number of steps are needed for the system to reach equilibrium, and the discrepancy in the size of the nanoparticles in Ref. 7 and Ref. 22 can be narrowed by increasing the simulation steps to mimic the whole heat-treatment process. We emphasize that the distributed implementation presented in this is the reason that we can run such a large-number of MC steps in a time budget of a few days.
Although there is no time dependence in MC method, the time scale of the simulation can be estimated using a simple diffusion model, similar to kinetic MC. A well-known model to describe the diffusion-controlled coarsening is Lifshitz–Slyozov–Wagner (LSW) theory, which gives
$r(t)^{3}-r(0)^{3}=Kt$ (9)
where $r(t)$ is the average precipitate radius at time t, and K is the diffusion speed, which depends exponentially on the energy barrier for swapping the atoms. Note that the experimental diameter of the NP is approximately 30 nm, and it takes about 2.9 M MC steps to grow from 7nm to 14 nm at T=1000 K, therefore, the number of MC steps needed to reach the experimental size is
$n_{\text{exp}}=\frac{30^{3}}{(14^{3}-7^{3})}\times\left(3\times 10^{6}-10^{5}\right)=3.26\times 10^{7}$ (10)
[…]
$t_{\text{sim}}=\frac{3600\times 3\times 10^{6}}{3.26\times 10^{7}}s=331\;s$ (11)
Of course, this is a very crude estimation, but it indicate that our MC simulation indeed reaches the time regime of real physical process of heat-treatment.
4 Conclusions
In this work, we have presented a distributed framework for the SMC-X method, demonstrating its versatility with implementations on both CPU and GPU architectures. By leveraging hierarchical parallelism, data-locality optimization, and computation–communication overlap, the distributed SMC-X achieves exceptional performance in system size, simulation throughput, and scaling efficiency. When combined with DFT-trained machine learning models, SMC-X attains first-principles accuracy while extending both the spatial and temporal scales of atomistic simulations to previously inaccessible regimes. Such capability is essential for predicting large-scale chemical defects in chemically complex alloys and, in turn, for elucidating the origin of their exceptional properties. This claim is exemplified by our billion-atom simulations of FeCoNiAlTi, where the extended spatial and temporal scales prove critical for bridging the gap between experimental observations and theoretical predictions. Altogether, our results demonstrate that the SMC-X + ML framework sheds light on the challenging yet critical problem of simulation-driven design of chemically complex materials.
Data Availability
All the dadta data will be publically available in Zenodo upon the acceptance of this paper.
Code Availability
A binary version of the code is available upon reasonable request.
Author contributions
X.L. conceived and supervised the project, developed the code prototype, fitted the ML model, and wrote the original manuscript. K.Y. implemented the production code and collected the original data. F.Z. performed data analysis and visualization. P.X. acquired and managed the computing resources. All authors discussed the results and contributed to the final version of the manuscript.
{acknowledgement}
The work of X. Liu and F. Zhou was supported by the National Natural Science Foundation of China under Grant 12404283. X. Liu acknowledges helpful discussion with Zongrui Pei on precipitate sizes in alloy.
Competing Interests
The authors declare no competing interests.

### v2

@@
 In this work, we advance the recently proposed SMC-X method through distributed computation on either GPUs or CPUs, pushing both spatial and temporal scales of atomistic simulation of chemically complex alloys to previously inaccessible scales.
-This includes a 16-billion-atom HEA system extending to the micrometer regime in space, and a 1-billion-atom HEA evolved over more than three million Monte Carlo swap steps, approaching the minute regime in time.
+This includes a record-breaking 128-billion-atom HEA system extending to the micrometer regime in space, and a 1-billion-atom HEA evolved over more than three million Monte Carlo swap steps, approaching the minute regime in time.
 We show that such large-scale simulations are essential for bridging the gap between experimental observations and theoretical predictions of the nanoprecipitate sizes in HEAs, based on analysis using the Lifshitz–Slyozov–Wagner (LSW) theory for diffusion-controlled coarsening.
@@
 However, the vast chemical complexity present in many technologically important materials poses formidable challenges for existing computational approaches.
-This challenge is particularly pronounced in high-entropy alloys (HEAs) 1, which contain multiple principal elements and therefore exhibit substantially increased chemical complexity.
+This challenge is particularly pronounced in high-entropy alloys (HEAs) EGeorge_Nature, which contain multiple principal elements and therefore exhibit substantially increased chemical complexity.
 Moreover, this complexity is further compounded by the interplay and competition between order and disorder, as well as enthalpic and entropic effects, across diverse spatial, temporal, and temperature scales, making predictive modeling of such systems exceptionally demanding.
-Capturing these complexities, including chemical composition 2, disorder 3, chemical short-range order (SRO) 4, 5, chemical medium-range order (MRO) 6, multi-component intermetallic nanoparticles (MCINP) 7, and nano lamellae8, 9, requires modeling and simulating atomic systems over large spatial and temporal scales, with accuracy and generalizability of first principles method 10, 11, 12.
-In recent years, machine learning potentials (MLPs) 13, 14, 15 trained on DFT datasets 16, 17 have emerged as a highly promising solution to the aforementioned bottlenecks 18, 19.
-By fitting interatomic potential energy surfaces with machine learning models, MLPs can retain near-DFT accuracy 20 while dramatically increasing simulation efficiency and extending accessible system sizes 21.
-With an appropriate interaction cut-off radius, many MLPs can reduce the computational complexity of traditional DFT calculations from cubic scaling, $O(N^{3})$, to linear scaling, $O(N)$ 15, with respect to the number of atoms $N$.
-Moreover, by bypassing the explicit solution of the electronic structure, the computational prefactor can be reduced by orders of magnitude 22, enabling simulations at spatial scales previously deemed intractable 23.
-The integration of MLPs with molecular dynamics (MD), a foundational atomistic simulation technique renowned for its excellent scaling, has revolutionized computational materials science 24, chemistry 25, and biology 26 by enabling accurate, large-scale simulations at significantly reduced cost.
-Although the MLP + MD approach successfully scales simulations to unprecedented sizes 27, 28, 29, its intrinsic limitations restrict accessible time scales to the nanosecond regime, which typically requires millions of sequential MD steps.
-Consequently, it is extremely challenging, if not impossible, for MD to capture many important slow-evolving phenomena, such as precipitate formation in alloys 7.
+Capturing these complexities, including chemical composition ChemicalComposition2019Nature, disorder AtomComplexity, chemical short-range order (SRO) SRO_NCS_2023, APT_CoCrNi_NatureMat_2024, chemical medium-range order (MRO) mediumRangeOrderNC, multi-component intermetallic nanoparticles (MCINP) Yang933, and nano lamellaeCTLiuNC, MaxPlankNC2025, requires modeling and simulating atomic systems over large spatial and temporal scales, with accuracy and generalizability of first principles method LIU2023101018, LIU2019107955, Eisenbach_2019.
+In recent years, machine learning potentials (MLPs) JCTCBP, Zhang_2025, MPNNs_2023 trained on DFT datasets OQMD, Horton2025 have emerged as a highly promising solution to the aforementioned bottlenecks UniversalModel_NCS, 2025Nature.
+By fitting interatomic potential energy surfaces with machine learning models, MLPs can retain near-DFT accuracy Ong_JPC_2020 while dramatically increasing simulation efficiency and extending accessible system sizes doi:10.1126/science.abn3445.
+With an appropriate interaction cut-off radius, many MLPs can reduce the computational complexity of traditional DFT calculations from cubic scaling, $O(N^{3})$, to linear scaling, $O(N)$ MPNNs_2023, with respect to the number of atoms $N$.
+Moreover, by bypassing the explicit solution of the electronic structure, the computational prefactor can be reduced by orders of magnitude liuNPJ2025, enabling simulations at spatial scales previously deemed intractable 10.1145/3458817.3487400.
+The integration of MLPs with molecular dynamics (MD), a foundational atomistic simulation technique renowned for its excellent scaling, has revolutionized computational materials science 2025Nature, chemistry Zhang2024, and biology Gkeka2020 by enabling accurate, large-scale simulations at significantly reduced cost.
+Although the MLP + MD approach successfully scales simulations to unprecedented sizes 10.1145/3581784.3627041, 2020NPJ_Li, Yin2021, its intrinsic limitations restrict accessible time scales to the nanosecond regime, which typically requires millions of sequential MD steps.
+Consequently, it is extremely challenging, if not impossible, for MD to capture many important slow-evolving phenomena, such as precipitate formation in alloys Yang933.
 For instance, a single swap step between two neighboring lattice sites may take millions or even billions of MD steps to simulate.
-For such a purpose, Monte Carlo (MC) simulations offer unique advantages due to their characteristic of directly sampling configurational space, rather than tracking real-time trajectories as in MD 30, 31.
-However, compared to MD, the integration of conventional MC with MLP faces difficulties when scaled up to large systems 22.
-This challenge stems from the intrinsically sequential update nature in mainstream Monte Carlo algorithms, such as the Metropolis method, where trial moves are proposed site by site, inherently limiting parallelization 32.
-This bottleneck is in stark contrast with MD, where all atoms can be updated simultaneously at each step.
-Previous efforts to scale Monte Carlo (MC) simulations have introduced parallel trial moves using generalized checkerboard algorithms, with most implementations customized for specific interaction types or hardwares 33, 34, 35, 36, 37, 32, and only a few designed for general interactions 38, 22.
-Among all these implementations, the recently proposed scalable Monte Carlo at eXtreme (SMC-X) 22 method stands out due to its superb efficiency.
+For such a purpose, Monte Carlo (MC) simulations offer unique advantages due to their characteristic of directly sampling configurational space, rather than tracking real-time trajectories as in MD Korman_npj, LIU2021110135.
+However, compared to MD, the integration of conventional MC with MLP faces difficulties when scaled up to large systems liuNPJ2025.
+This challenge stems from the intrinsically sequential update nature in mainstream Monte Carlo algorithms, such as the Metropolis method, where trial moves are proposed site by site, inherently limiting parallelization Anderson2016SMC.
+This bottleneck is in stark contrast to that in MD, where all atoms can be updated simultaneously at each step.
+Previous efforts to scale Monte Carlo (MC) simulations have introduced parallel trial moves using generalized checkerboard algorithms, with most implementations customized for specific interaction types or hardware PREIS20094468, BLOCK20101549, ROMERO2020107473, liu2019high, OrtegaZamorano2013FPGAIsing, Anderson2016SMC, and only a few designed for general interactions sadigh2012scalable, liuNPJ2025.
+Among all these implementations, the recently proposed scalable Monte Carlo at eXtreme (SMC-X) liuNPJ2025 method stands out due to its superb efficiency.
 Such a high efficiency is achieved via exposing the previously hidden degree of parallelism by dynamical link-cell (LC) decomposition and local interaction zone (LIZ), and by carefully mapping these degrees of parallelisms onto modern GPU architectures to remove unnecessary data movement.
-These innovations allow SMC-X to simulate systems as large as one billion atoms using a single GPU 22.
-This performance not only substantially exceeds all prior MC + MLP methods, but also exhibits advantages against state-of-the-art (SOTA) MD + MLP implementations in both the number of atoms and the atomistic simulation throughput, measured by $N_{\rm atom}\times N_{\rm steps}$ per second 22.
+These innovations allow SMC-X to simulate systems as large as one billion atoms using a single GPU liuNPJ2025.
+This performance not only substantially exceeds all prior MC + MLP methods, but also exhibits advantages against state-of-the-art (SOTA) MD + MLP implementations in both the number of atoms and the atomistic simulation throughput, measured by $N_{\rm atom}\times N_{\rm steps}$ per second liuNPJ2025.
 A comparison of various computational methods for chemically complex alloys is shown in Tab.
 1.
-It should be noted that while mesoscale methods such as the Potts model, kinetic Monte Carlo, and phase-fild methods can simulate very large spatial and temporal scales, they introduce coarse-grained approximations that rely on empirical parameters, making them less rigorous at the atomistic level, therefore are to compared in Tab.
+It should be noted that while mesoscale methods such as the Potts model, kinetic Monte Carlo, and phase-fild methods can simulate very large spatial and temporal scales, they introduce coarse-grained approximations that rely on empirical parameters, making them less rigorous at the atomistic level, therefore are not listed in Tab.
 1.
@@
 Note that the basic concept of the SMC-X algorithm has been introduced in Ref.
-22, but the focus was on its application in materials science.
+liuNPJ2025, but the focus was on its application in materials science.
 Here we present a more detailed discussion and introduce the distributed implementation on both CPUs and GPUs.
@@
 Physical observables $\langle\hat{O}\rangle$ such as the expected energy, magnetization, specific heat, and magnetic susceptibility can then be calculated accordingly from
-$\langle\hat{O}\rangle=\sum_{k}\frac{O(\sigma_{k})e^{-\beta H(\bm{\sigma}_{k})}}{Z}.$ (2)
+$\langle\hat{O}\rangle=\sum_{k}\frac{O(\sigma_{k})e^{-\beta H(\bm{\sigma}_{k})}}{e^{-\beta H(\bm{\sigma}_{k})}}.$ (2)
 A sample following the probability distribution of Eq.
-1 can be generated efficiently via Markov chain Monte Carlo (MCMC) simulation.
+1 can be efficiently generated via a Markov chain Monte Carlo (MCMC) simulation.
 For such a purpose, the Metropolis-Hastings algorithms is the most commonly used method.
@@
 The trial moves within the Metropolis-Hasting algorithm are inherently sequential, which poses a significant challenge to parallelization.
-Although advanced techniques such as cluster MC 39 or the Wang–Landau algorithms 40, 41, 42 can alleviate this limitation by reducing or avoiding sequential site-by-site updating, they are often restricted in scope and lack general applicability.
+Although advanced techniques such as cluster MC PhysRevLett.58.86 or the Wang–Landau algorithms PhysRevLett.86.2050, 10.1145/1654059.1654125, PhysRevE.97.043301 can alleviate this limitation by reducing or avoiding sequential site-by-site updating, they are often restricted in scope and lack general applicability.
 A more general approach leverages the assumption of short-range interactions: If interactions are sufficiently localized, then spatially distant regions can be updated independently and in parallel.
 In the simple Ising model, this principle underlies the well-known checkerboard algorithm.
-Extensions of this approach to three-dimensional lattices and beyond nearest-neighbor interactions are conceptually straightforward for specific models 33.
+Extensions of this approach to three-dimensional lattices and beyond nearest-neighbor interactions are conceptually straightforward for specific models PREIS20094468.
 However, generalization to arbitrary short-range interactions is more limited.
-One notable example is the Scalable Parallel Monte Carlo (SPMC) algorithm 38, as implemented in the LAMMPS package 43.
-Despite its utility, SPMC has disadvantage for further scaling up: to minimize the weak spatial correlations in SPMC algorithm, the total volume of the domains need to be maximized 38, which underutilizes the available parallelism, especially on architectures with a high degree of concurrency.
-As a result, current application of SPMC are limited systems of less than a million atoms 29, 44.
+One notable example is the Scalable Parallel Monte Carlo (SPMC) algorithm sadigh2012scalable, as implemented in the LAMMPS package THOMPSON2022108171.
+Despite its utility, SPMC has disadvantage for further scaling-up: to minimize the weak spatial correlations in SPMC algorithm, the total volume of the domains need to be maximized sadigh2012scalable, which underutilizes the available parallelism, especially on architectures with a high degree of concurrency.
+As a result, current applications of SPMC are limited systems of less than a million atoms Yin2021, YAO2024120457.
 Figure 2: A schematic to illustrate the SMC-X algorithm, including link-cell (LC), local interaction zone (LIZ), and domain decomposition scheme, using the 2D square lattice as an example.
@@
 Adapted from Ref.
-22
+liuNPJ2025
 2.2 Distributed SMC-X
-In the SMC-X framework 22, the challenge of site-by-site sequential updating is addressed through two key innovations: dynamical link-cell (LC) and local interaction zone (LIZ), as illustrated in Fig.
+In the SMC-X framework liuNPJ2025, the challenge of site-by-site sequential updating is addressed through two key innovations: dynamical link-cell (LC) and local interaction zone (LIZ), as illustrated in Fig.
 2.
-Compared to SPMC 38, the LCs are dynamically constructed to ensure that every site is visited exactly once per MC step, and that the detailed balance condition is strictly satisfied.
+Compared to SPMC sadigh2012scalable, the LCs are dynamically constructed to ensure that every site is visited exactly once per MC step, and that the detailed balance condition is strictly satisfied.
 The resulting degree of parallelism is given by:
@@
 By choosing the smallest LC size that fully separates short-range interactions, the parallelism $P_{\mathrm{LC}}$ can be maximized.
-The LIZ is introduced to harness the short-range interaction to reduce the computational complexity of a MC trial move from $O(N)$ to $O(1)$.
-Furthermore, by computing the atomic features within each LIZ on the fly and tightly couple it with the MC simulation, the memory usage in SMC-X can be drastically reduced, and near-core caches can be effectively utilized to significantly enhance the computational throughput.
+The LIZ is introduced to exploit the short-range interaction to reduce the computational complexity of an MC trial move from $O(N)$ to $O(1)$.
+Furthermore, by computing the atomic features within each LIZ on the fly and tightly coupling it with the MC simulation, the memory usage in SMC-X can be drastically reduced, and near-core caches can be effectively utilized to significantly enhance computational throughput.
 In addition to LC and LIZ, there are additional parallel opportunities in SMC-X, as illustrated in the algorithm part in Fig.
@@
 2.
-Under this scheme, if the number of atoms per GPU remains fixed, the total system size then increases proportionally to the number of GPUs, demonstrating weak scaling.
+Under this scheme, if the number of atoms per GPU remains fixed, the total system size increases proportionally to the number of GPUs, demonstrating weak scaling.
 Conversely, when the same system is distributed over more GPUs, the time required for a single MC step decreases, enabling more steps to be simulated within the same wall-clock time, an indication of strong scaling.
@@
 Therefore, the MC simulation is embarrassingly parallel when properties at different temperatures are needed, such as to search for an order-disorder transition temperature.
-On the other hand, it is well known that the convergence speed of conventional MC can be slow, particularly near second-order transition, and the replica-exchange (RE, also known as parallel tempering) is a widely used technique to address this problem.
+On the other hand, it is well known that the convergence speed of conventional MC can be slow, particularly near second-order transition, and replica-exchange (RE, also known as parallel tempering) is a widely used technique to address this problem.
 In RE, multiple replicas of the same system are simulated concurrently at different temperatures.
@@
 By exchanging configurations between neighboring temperatures, the replica-exchange algorithm mitigates the risk of the system becoming trapped in local energy minima.
-This is especially critical near second-order phase transition points 39, 40, where the divergence of the correlation length makes it challenging for simulations at a fixed temperature to escape metastable states.
+This is especially critical near second-order phase transition points PhysRevLett.58.86, PhysRevLett.86.2050, where the divergence of the correlation length makes it challenging for simulations at a fixed temperature to escape metastable states.
 For large systems, an additional complication needs to be considered: the energy fluctuation decreases as the system size increases.
@@
 Note that the $\rm{Fe_{29}Co_{29}Ni_{28}Al_{7}Ti_{7}}$ HEA studied in this work does not show obvious order-disorder transition in its $C_{v}$ curve, as illustrated in Ref.
-22.
+liuNPJ2025.
 2.3 SMC-GPU and SMC-CPU
@@
 Two models, EPI and qSRO, are used.
-The EPI model31, 45 contains only pairwise interactions, which is computationally highly efficient since there is no need to explicitly evaluate the local energy changes within the LIZ.
-On the other hand, the qSRO model 22 represents a general nonlinear short-range model, where the local energies of all lattice sites within LIZ are calculated to determine the energy change due to a MC move.
+The EPI modelLIU2021110135, ZHANG2020108247 contains only pairwise interactions, which is computationally highly efficient since there is no need to explicitly evaluate the local energy changes within the LIZ.
+On the other hand, the qSRO model liuNPJ2025 represents a general nonlinear short-range model, where the local energies of all the lattice sites within LIZ are calculated to determine the energy change due to a MC move.
 For SMC-GPU, the computational platform is a 2-node GPU cluster.
@@
 Each CPU contains multiple NUMA domains, with dozens of physical cores in each domain.
-The DFT dataset is calculated with the LSMS method 46.
+The DFT data-set is calculated with the LSMS method PhysRevLett.75.2867.
 LSMS is an all-electron electronic structure calculation method in which the computational cost scales linearly with respect to the number of atoms.
@@
 We used PBE as the exchange-correlational functional.
-To enhance the representativeness of the DFT data, the total dataset is made up of 120 randomly generated configurations and 100 configurations from Monte Carlo simulation, as proposed in 31.
+To enhance the representativeness of the DFT data, the total dataset is made up of 120 randomly generated configurations and 100 configurations from Monte Carlo simulation, as proposed in LIU2021110135.
 Therefore, the total number of configurations is 220, with each comprising 100 atoms.
 3.1 Performance
-Figure 4: Comparison of different ML accelerated atomistic simulation methods in terms of atom per chip and throughput per chip.
+(a) Atoms per chip comparison
+(b) Throughput per chip comparison
+(c) Overall performance metrics
+Figure 4: Comparison of different ML accelerated atomistic simulation methods: (a) shows the number of atoms that can be simulated per chip, (b) displays the computational throughput achievable per chip, and (c) comparison of the performance of our results with previous SOTA in terms of total $N_{atom}$ ($x$-axis), total throughput ($y$-axis), throughput per chip (color), and the number of chips (circles with radius proportional to $N_{Chip}^{1/3}$).
+See original data at Tab.
+2.
 Table 2: Comparison of the performance of atomistic simulations achieving DFT-level accuracy (10 meV/atom).
@@
 Simu. Potential Sys. Chips Machine Atoms TP TP/Chip
-LAMMPS 27 Allegro HIV 5,120 Perlmutter 100 M $1.27\times 10^{10}$ $2.47\times 10^{6}$}
-LAMMPS 47 DP H2O 90,000 Sunway 29 B $1.14\times 10^{10}$} $1.26\times 10^{5}$
-LAMMPS 48 TensorMD W 32,000 Sunway 51.8 B $\mathbf{{1.30\times 10^{10}}}$ $4.06\times 10^{5}$
-GPUMD 49 UNEP-v1 Cu 8 A100 1 node 100 M $1.50\times 10^{8}$} $1.86\times 10^{7}$
-GPUMD 50 NEP89 HEA 1 H800 1 node 15 M $7.0\times 10^{6}$} $7.0\times 10^{6}$
+LAMMPS 10.1145/3581784.3627041 Allegro HIV 5,120 A100 Perlmutter 100 M $1.27\times 10^{10}$ $2.47\times 10^{6}$}
+LAMMPS 10880101 DP H2O 90,000 CPU Sunway 29 B $1.14\times 10^{10}$} $1.26\times 10^{5}$
+LAMMPS 10.1145/3503221.3508425 DP H2O 27,360 V100 Summit 3.9B $1.12\times 10^{10}$ $4.11\times 10^{5}$
+LAMMPS 10.1145/3503221.3508425 DP H2O 9,936 CPU Fugaku 1.5B $2.17\times 10^{9}$ $2.18\times 10^{5}$
+LAMMPS 10.1145/3458817.3487400 SNAP C 27,900 V100 Summit 20B $\mathbf{2.89\times 10^{10}}$ $1.08\times 10^{6}$
+LAMMPS 10.1145/3710848.3710882 TensorMD W 32,000 CPU Sunway 51.8 B $1.30\times 10^{10}$}} $4.06\times 10^{5}$
+GPUMD MLPAlloys2024 UNEP-v1 Cu 8 A100 1 node 100 M $1.50\times 10^{8}$} $1.86\times 10^{7}$
+GPUMD liang2025nep89universalneuroevolutionpotential NEP89 HEA 1 H800 1 node 15 M $7.0\times 10^{6}$} $7.0\times 10^{6}$
 SMC-X EPI HEA 16 H800 2-node 16 B $\mathbf{5.63\times 10^{10}}$ $\mathbf{3.52\times 10^{9}}$
 SMC-X qSRO HEA 16 H800 2-node 16 B $5.79\times 10^{8}$ $\mathbf{3.62\times 10^{7}}$}
+SMC-X qSRO HEA 32 H800 4-node 128 B $1.17\times 10^{9}$ $3.29\times 10^{7}$
 SMC-X qSRO HEA 2 CPU 1 node 16 B $4.44\times 10^{7}$ $2.22\times 10^{7}$}}
-The performance is measured with three metrics: number of atoms, which measure the spatial scale; simulation throughput, which measures time to solution; and scaling efficiency, which measures the scale-up potential with increasing computational resources.
+Performance is measured with three metrics: number of atoms, which measure spatial scale; simulation throughput, which measures time to solution; and scaling efficiency, which measures the scaling potential with increasing computational resources.
 As shown in Fig.
-4, the EPI model exhibits exceptional computational efficiency on a single GPU, achieving a throughput of $3.5\times 10^{9}$ $\text{atom}\cdot\text{step}/\text{s}$—two orders of magnitude faster than the qSRO_GPU result of $3.7\times 10^{7}$ $\text{atom}\cdot\text{step}/\text{s}$.
+6(b), the EPI model exhibits exceptional computational efficiency on a single GPU, achieving a throughput of $3.5\times 10^{9}$ $\text{atom}\cdot\text{step}/\text{s}$—two orders of magnitude faster than the qSRO_GPU result of $3.7\times 10^{7}$ $\text{atom}\cdot\text{step}/\text{s}$.
 In comparison, the single-NUMA performance of qSRO_CPU reaches $3.5\times 10^{6}$ $\text{atom}\cdot\text{step}/\text{s}$, or $2.5\times 10^{7}$ $\text{atom}\cdot\text{step}/\text{s}$ per CPU.
-The throughput of SMC-X not only far exceeds that of traditional Monte Carlo methods but also outperforms state-of-the-art molecular dynamics implementations, as shown in Fig.
+The throughput of SMC-X not only far exceeds that of traditional Monte Carlo methods, but also outperforms state-of-the-art molecular dynamics implementations, as shown in Fig.
 4 and summarized in Tab.
 2.
-The advantage is even more pronounced in terms of the number of atoms supported on a single chip: the SMC-GPU implementation accommodates up to 1 billion atoms on a single GPU, whereas the best reported MD simulation to date reaches only 15 million atoms per GPU, as shown in the row of NEP89 in Tab.
+It is noteworthy that SMC-X, utilizing merely 32 H800 GPUs, can simulate systems of 128 billion atoms, which is a scale that surpasses all other ML-accelerated MC or MD simulations, including those performed on some of the world’s most powerful supercomputers, as shown in Fig.
+4(c).
+The advantage is even more pronounced in terms of the number of atoms supported on a single chip: the SMC-GPU implementation accommodates up to 4 billion atoms on a single GPU by running 4 MPI ranks simultaneously, whereas the best reported MD simulation to date reaches only 15 million atoms per GPU, as shown in the row of NEP89 in Tab.
 2.
-As mentioned, these performance gains stem from key implementation optimizations, including on-the-fly computation and tight coupling between the atomistic simulation and machine learning models, which collectively minimize data movement and maximize computational efficiency.
+As mentioned, these performance gains stem from key implementation optimizations, including on-the-fly computation and tight coupling between atomistic simulation and machine learning models, which collectively minimize data movement and maximize computational efficiency.
 Figure 5: The strong and weak scaling of SMC-GPU and SMC-CPU with EPI and qSRO model. CCO represents computation-communication-overlap.
@@
 Finally, although both strong and weak scaling efficiencies for qSRO_CPU are generally good, they are consistently lower than their GPU counterparts, indicating that cross-CPU NUMA communication is more costly than GPU interconnects such as NVLINK.
+Finally, we comment on the limitations of current SMC-X implementations and its future potential.
+The present implementation requires kernel-level programming to optimize the machine-learning (ML) potential, which enables peak performance but makes the integration of additional potentials labor-intensive.
+Consequently, we have so far implemented only a simple qSRO model as a representative of general short-range models.
+This naturally raises the question of whether such performance can be sustained with more complex ML interatomic potentials (MLIPs).
+To investigate this, we performed a roofline analysis of the SMC-GPU code using NVIDIA Nsight tools.
+The results show that the current implementation is bounded primarily by integer arithmetic for on-the-fly index calculations and by HBM bandwidth for data movement, whereas the floating-point units, including CUDA vector units and Tensor Cores, remain significantly underutilized due to the low theoretical arithmetic intensity of about 1.
+Given that modern GPUs are designed for workloads with arithmetic intensities in the hundreds or higher, it is reasonable to expect that incorporating more complex models with higher computational intensity will not compromise performance.
+Indeed, in a prototype implementation using a three-layer multilayer perceptron (MLP) with a hidden dimension of 256, we observe that the excellent speed reported for qSRO is retained.
+Looking ahead, accommodating fully pretrained MLIPs may require a modest redesign of the SMC-X code architecture, potentially introducing a small performance overhead.
+Nevertheless, the method would remain exceptionally fast compared with conventional Monte Carlo approaches, as the dominant speedup arises from the SMC-X algorithm itself.
+These enhancements will be explored in future work.
 3.2 Evolution of chemical defects
-The Warren–Cowley short-range order (SRO) parameter 51, 31, denoted as $\alpha_{m}^{pp^{\prime}}$, is widely used to quantify local chemical ordering. It is defined as
+The Warren–Cowley short-range order (SRO) parameter PhysRev.77.669, LIU2021110135, denoted $\alpha_{m}^{pp^{\prime}}$, is widely used to quantify local chemical ordering. It is defined as
 $\alpha_{m}^{pp^{\prime}}=1-\frac{P^{p|p^{\prime}}_{m}}{c_{p}},$ (8)
@@
 In particular, Ti and Ni atoms display a strong tendency to segregate from Fe and Co, with the Ti–Fe pair showing the strongest separation.
-This behavior is consistent with previous experimental 7 and theoretical 22 studies, which reported the formation of $L1_{2}$-type nanoprecipitates enriched in Ni and Ti within a Fe- and Co-rich matrix.
-Figure 6: The short range order parameters calculated with the 1 million atom supercell. (a) The SROs for different elements pairs; (b) The SROs for the same element pairs.
+This behavior is consistent with previous experimental Yang933 and theoretical liuNPJ2025 studies, which reported the formation of $L1_{2}$-type nanoprecipitates enriched in Ni and Ti within a Fe- and Co-rich matrix.
+(a) SRO of different element pairs
+(b) SRO of identical element pairs
+Figure 6: The chemical short-range order parameters (SROs) calculated with the 1 million atom supercell, estimated with 9 configuration samples.
+The error bars are estimated with the difference between the largest and smallest ones among the 9 SRO values.(a) The SROs for different elements pairs; (b) The SROs for the same element pairs.
 To investigate microstructure evolution, we performed large-scale Monte Carlo simulations using the EPI model, with a supercell comprising 1,066,867,200 atoms (dimensions: $672\times 630\times 630\times 4$).
-The computing platform is the two node, 16 H800 GPU cluster, and the ML model is the EPI model to maximize the simulation speed.
-We decrease the temperature from 2000 K to 1000 K, with tens of thousands of MC steps at each temperature, except for a temperature we refer to as the “aging” temperature $T_{aging}$, at which the simulation run for 3 million MC steps .
+The computing platform is the two-node, 16-H800 GPU cluster, and the ML model is the EPI model to maximize the simulation speed.
+We decrease the temperature from 2000 K to 1000 K, with tens of thousands of MC steps at each temperature, except for a temperature we refer to as the “aging” temperature $T_{aging}$, at which the simulation runs for 3 million MC steps .
 The X-Y face of the system for $T_{aging}=1000$ K experiment is shown in Fig.
-7 (a, b), where we used 8 GPUs for lattice parallelism and 2 temperature degree of parallelism, with $\Delta T=0.06$ K.
-As mentioned in the method section, such a small $\Delta T$ is for reasonable acceptance ratio of RE in large supercells.
+7 (a, b), where we used 8 GPUs for lattice parallelism and 2 temperature degrees of parallelism, with $\Delta T=0.06$ K.
+As mentioned in the Methods section, such a small $\Delta T$ is for a reasonable acceptance ratio of RE in large supercells.
 The results for $T_{aging}=1500$ K is shown in Fig.
 7 (c, d), where no RE is used, and the lattice is parallelized over the total 16 GPUs.
-From these figures, it can be seen that the NPs rich in Ni and Ti already forms at 1500 K.
+From these figures, it can be seen that the NPs rich in Ni and Ti already form at 1500 K.
 Upon further cooling, nanoprecipitates gradually grow in size.
@@
 (a) The APT sample extracted from SMC-X simulation in Ref.
-22.
-(b) The APT reconstructed sample image of the same materials from experiment 7, where MCM represents multicomponent matrix, and MCINP represents multicomponent intermetallic nanoparticle.
+liuNPJ2025.
+Reproduced from Ref.
+liuNPJ2025.
+Available under a $C$-BY NC-ND 4.0 license.
+Copyright Xianglin Liu et al.
+(b) The APT reconstructed sample image of the same materials from experiment Yang933, where MCM represents multicomponent matrix, and MCINP represents multicomponent intermetallic nanoparticle.
+Reproduced with permission from Ref.
+Yang933.
+Copyright 2018 AAAS.
 (c) The APT sample from this work, using 3 million MC steps at 1000 K.
@@
 (a-c) blue represents MCINP, and red represents MCM.
-To reveal further details of the NPs, we extract an atom-probe-tomography (APT) sample from the simulation results and compared them with previous experimental and theoretical studies, as shown in Fig.
+(e) Averaged compositions obtained from 9 nearby cylinderical specimen of r=12ar=12\;a and l=72al=72\;a, where $a$ is the lattice constant.
+To reveal further details of the NPs, we extract an atom-probe-tomography (APT) sample from the simulation results and compare them with previous experimental and theoretical studies, as shown in Fig.
 8.
 First, we see that the previous study shown in Fig.
-8(a) clearly underestimates the sizes of the NPs, which is about 7 nm in diameter, as compared to the experimental result in Fig.
+8(a) clearly underestimates the sizes of the NPs, which is about 7 nm in diameter, compared to the experimental result in Fig.
 8 (b), which is about 30 nm.
-Second, we see that this gap between experiment and theory can be reduced by increasing the MC steps from 0.1 million to 3 million, which enhances the size of the NP to about 14 nm.
+Second, we see that this gap between experiment and theory can be reduced by increasing the MC steps from 0.1 million to 3 million, which increases the size of the NP to about 14 nm.
 Finally, note that the experimental sample is synthesized by homogenizing at 1423 K for 2 h and aged at 1053 K for 4 h.
@@
 It can be seen that the size of the NP can be further increased to about 20 nm.
-These observations reveal that larger number of steps are needed for the system to reach equilibrium, and the discrepancy in the size of the nanoparticles in Ref.
-7 and Ref.
-22 can be narrowed by increasing the simulation steps to mimic the whole heat-treatment process.
+These observations reveal that a larger number of steps are needed for the system to reach equilibrium, and the discrepancy in the size of the nanoparticles in Ref.
+Yang933 and Ref.
+liuNPJ2025 can be narrowed by increasing the simulation steps to mimic the whole heat-treatment process.
 We emphasize that the distributed implementation presented in this is the reason that we can run such a large-number of MC steps in a time budget of a few days.
-Although there is no time dependence in MC method, the time scale of the simulation can be estimated using a simple diffusion model, similar to kinetic MC.
+The chemical composition at 1000 K is analyzed, as shown in Fig 8 (e).
+The composition profile is extracted from 9 nearby cylindrical specimen of radius r=12ar=12\;a and length of l=72al=72\;a, where $a$ is the lattice constant.
+It can be seen that Fe and Co are favored in the disordered matrix phase, while Ni and Ti and strongly favored in the ordered (Ni,Co)3-(Ti,Al) $L1_{2}$ phase.
+These results agrees well with both experiments Yang933 and previous theoretical studies liuNPJ2025, WANG2025120635.
+The error bar is the standard deviation of the compositions.
+Although there is no time dependence in the MC method, the time scale of the simulation can be estimated using a simple diffusion model, similar to kinetic MC.
 A well-known model to describe the diffusion-controlled coarsening is Lifshitz–Slyozov–Wagner (LSW) theory, which gives
@@
 $t_{\text{sim}}=\frac{3600\times 3\times 10^{6}}{3.26\times 10^{7}}s=331\;s$ (11)
-Of course, this is a very crude estimation, but it indicate that our MC simulation indeed reaches the time regime of real physical process of heat-treatment.
+Of course, this is a very crude estimation, but it indicates that our MC simulation indeed reaches the time regime of a real physical process of heat-treatment.
 4 Conclusions
 In this work, we have presented a distributed framework for the SMC-X method, demonstrating its versatility with implementations on both CPU and GPU architectures.
-By leveraging hierarchical parallelism, data-locality optimization, and computation–communication overlap, the distributed SMC-X achieves exceptional performance in system size, simulation throughput, and scaling efficiency.
+Using hierarchical parallelism, data-locality optimization, and computation–communication overlap, the distributed SMC-X achieves exceptional performance in simulation throughput and scaling efficiency.
+In particular, it achieves the largest system size for ML-accelerated atomistic simulation, reaching 128 billion atoms using only 32 H800 GPUs, a much smaller hardware budget compared to previous results obtained with the largest supercomputers.
 When combined with DFT-trained machine learning models, SMC-X attains first-principles accuracy while extending both the spatial and temporal scales of atomistic simulations to previously inaccessible regimes.
 Such capability is essential for predicting large-scale chemical defects in chemically complex alloys and, in turn, for elucidating the origin of their exceptional properties.
-This claim is exemplified by our billion-atom simulations of FeCoNiAlTi, where the extended spatial and temporal scales prove critical for bridging the gap between experimental observations and theoretical predictions.
-Altogether, our results demonstrate that the SMC-X + ML framework sheds light on the challenging yet critical problem of simulation-driven design of chemically complex materials.
+This claim is exemplified by our billion-atom simulations of FeCoNiAlTi, where extended spatial and temporal scales are critical for bridging the gap between experimental observations and theoretical predictions.
+Together, our results demonstrate that the SMC-X + ML framework sheds light on the challenging but critical problem of simulation-driven design of chemically complex materials.
 Data Availability
-All the dadta data will be publically available in Zenodo upon the acceptance of this paper.
+[…]
 Code Availability
@@
 Zhou was supported by the National Natural Science Foundation of China under Grant 12404283.
+The work of F.
+Zhou was also supported by the High-Level Talent Start-up Fund provided by Xiangnan University.
 X.
@@
 Competing Interests
-The authors declare no competing interests.
+The authors declare no conflict of interest.